OpenStudy (anonymous):

In ∆ABC, AC = 15 centimeters, m B = 68°, and m C = 24°. What is BC to two decimal places?

OpenStudy (yolomcswagginsggg):

In a triangleABC, A+B+C=180 A=180-(B+C)=180-(68+42)=180-110=70 also sinA/BC=sinB/AC sin 70/BC=sin 68/15 BC=15*(sin 70/sin 68)=15*(0.939/0.927)=15*1.013=15.195